Winter Work Packet 2018 Th is packet is designed to help you study on kinematics and rotational motion during the break. It is requested that you complete this packet if you currently have less than a 65%, but all students are welcome to complete it. This packet will be due the first day of school, I will not accept it any later. I will be available to help with any questions regarding the problems via REMIND on Wednesday the second from 2PM to 5PM Section 1: Definitions and conversions 1. Perform the following conversions a. 30° to radians b. 2t radians to degrees c. 4 revolutions to radians d. 210° to radians When dealing with a rolling ball, which equation will tell us the linear distance the ball has travelled if we know the angle which it has covered? Show that by dividing by time, the conversion between tangential distance and 2. 3. ular distance becomes the conversion between tangential velocity and ang angular velocity What will happen to the following values if I were to double my radius: 4. a. Tangential velocity b. Angular velocity c. Centripetal force d. Moment of inertia e. Torque 5. Identify the following variables and which angular value they correspond to: a. b. C, θ e. r Section 2: Rotational CalculationsSTAF 6. I have a 2kg ball with a radius of.7m a. What is the moment of inertia of the ball? b. If I apply a force of 5 N to the edge of the ball at an angle of 80 the torque that is applied? c. For part c. What is the angular acceleration this will result in? d. If this acceleration continues for 10 seconds, what will the ve locity of the ball be? linearly? half hangs off the edge. e. If this acceleration continues for 10 seconds, how far will the ball roll 7. A meter stick is resting such that half of the ruler sits on a table, and the other a. If I apply a 50 g mass to the ruler at 20 cm from the edge of the ruler, and off the end of the table, what torque does this result in? b. What is the center of mass of the ruler-mass combo? c. Will this set up be stable? d. Repeat parts b. and c. for the ruler sitting with only 4 off the edge. 8. A flywheel is spinning at 4 revolutions per second a. If I apply a angular acceleration of Tt/3 rad/sec?, how long will it take to reach 8 revolutions per second? b. What angle will this acceleration result in measured in degrees? C. If the flywheel has a radius of .1 25m, what is the required centripetal acceleration? Section 2: Rotational Conceptual Questions 9. Considering what we have learned about Moment of Inertia, Center of Mass, 0. When you are in the front passenger seat of a car, and the driver makes a left and/or Torque, explain why a tightrope walker often carries a long pole. turn, you will often feel yourself pressed into the right side door. Why are you pressed into the right door? Why does the door press against you? Does your response make sense with Newtons laws? Does your response make sense Centripetal force?
Answer #1

1. To perform the following conversions which are given below as -

(a) 300 to radians

converting deg into rad :

\theta = [300 x (\pi / 1800)] rad

\theta = (\pi / 6) rad

\theta = 0.523 rad

(b) 2\pi radians to degree

converting rad to deg :

\theta = [2\pi x (1800 / \pi)] deg

\theta = 3600

(c) 4 revolutions to radians

converting rev to rad :

\theta = [4 x (2\pi)] rad

\theta = (8\pi) rad

\theta = 25.1 rad

(d) 2100 to radians

converting deg into rad :

\theta = [2100 x (\pi / 1800)] rad

\theta = (7\pi / 6) rad

\theta = 3.66 rad

5. To identify the following variables which are given below as -

(a) I \Rightarrow moment of inertia

(b) \omega\Rightarrow angular velocity

(c) \theta\Rightarrow angular distance

(d) \alpha\Rightarrow angular acceleration

(e) \tau\Rightarrow torque

Which angular value they correspond to :

\tau = I \alpha

6. (a) The moment of inertia of a ball which will be given by -

I = (2/5) m r2

I = [(0.4) (2 kg) (0.7 m)2]

I = 0.392 kg.m2

(b) The applied torque which will be given by -

\tau = r F sin \theta

\tau = (0.7 m) (5 N) sin 800

\tau = 3.44 N.m

(c) The angular acceleration which will be given by -

\alpha = \tau / I

\alpha = [(3.44 N.m) / (0.392 kg.m2)]

\alpha = 8.77 rad/s2

8. (a) Using equation of rotational motion (1), we have

\omegaf = \omegai + \alpha t

where, \omegai = initial angular velocity of a flywheel = 4 rev/s = 25.1 rad/s

\omegaf = final angular velocity of a flywheel = 8 rev/s = 50.2 rad/s

\alpha = angular acceleration = \pi/3 rad/s2 = 1.04 rad/s2

then, we get

t = [(50.2 rad/s) - (25.1 rad/s)] / (1.04 rad/s2)

t = [(25.1 rad/s) / (1.04 rad/s2)]

t = 24.1 sec

(b) Using equation of rotational motion (3), we have

\omegaf2 = \omegai2 + 2 \alpha\theta

(50.2 rad/s)2 = (25.1 rad/s)2 + 2 (1.04 rad/s2) \theta

[(2520.04 rad2/s2) - (630.01 rad2/s2)] = (2.08 rad/s2) \theta

\theta = [(1890.03 rad2/s2) / (2.08 rad/s2)]

\theta = 908.6 rad